Evaluate the function.
Given that f(x) = sqrt 6-x , find f(2.9).
A) sqrt 3.1 ≈ 1.55
B) sqrt 6 ≈ 2.4495
C) sqrt 3.1 ≈ 1.7607
D) sqrt 8.9 ≈ 2.9833

Answer by Alan3354(69443)

Given that f(x) = sqrt 6-x , find f(2.9).
------------------
Apparently you mean sqrt(6-x), not sqrt(6) - x. Use parentheses.
f(2.9) = sqrt(6-2.9) = sqrt(3.1)
= ~ 1.76
--------
A is also sqrt(3.1), but gives ~ 1.55 ???